OpenStudy (anonymous):

solve 2(x-11)-(x+6)=9x

OpenStudy (anonymous):

8x = -28 x = -28/8 x = -7/2

OpenStudy (anonymous):

can u show me how u got ur answer

OpenStudy (anonymous):

Expand the expression: 2(x-11)-(x+6) = 9x => 2x - 22 - x - 6 = 9x => x - 28 = 9x => 9x - x = -28 => 8x = -28 => x = -28/8 => x = -7/2

OpenStudy (anonymous):

thank you so much can i ask another

OpenStudy (anonymous):

Post as separate question. And, give a medal. :)